OpenStudy (anonymous):

How do you solve 1/3=x/30?

OpenStudy (anonymous):

you would cross-multiply: 1/3=x/30 1x=3(30) x=90

OpenStudy (anonymous):

i typed it wrong... hold on

OpenStudy (anonymous):

you would cross-multiply: 1/3=x/30 1(30)=3x 10=x

OpenStudy (anonymous):

there is an easier way... 1 x --- = --- 3 30 the number diagonal from the x is what you divide by the other two numbers you multiply (30 x 1)/3 = 10

OpenStudy (anonymous):

Thanks so much! I thought that was how to do it but it's been decades since I did this. My little boy needed help. :)

OpenStudy (anonymous):

It is good that you are helping him. :)